OpenStudy (anonymous):

if f(x)=3x^2-x, find f(m-4) what is the answer and can you explain how to work the problem

OpenStudy (anonymous):

f(m-4) = 3(m-4)^2 - (m-4) = 3(m^2 - 8m + 16) - m + 4 = 3m^2 - 25m + 52